The price elasticity of demand for the output of a profit-maximizing firm is E = −4. This firm will mark up the price of its product above marginal cost by __________ percent.

  • 25

  • 150

  • 50

  • 100

  • None of the options.
Answer #1

Answer


The correct answer is "option 1"
25

(P-MC)/P=-1/Ep

Ep=-4

(P-MC)/P=-1/-4

(P-MC)/P=1/4

(P-MC)/P=0.25 or 25%